RC meeting and Run Plan, Monday February 20th
From SBS wiki
Contents
RC Report
- https://logbooks.jlab.org/entry/4135224
- RC: Bogdan Wojtsekhowski, bogdanw@jlab.org
Summary
- Beam recovery was completed by 5am today
- We had ~45 uA beam until BLA for hall D took beam off
- Total charge collected over last seven shifts is just 1.7C due to the fire event and accelerator problems
Issues
- DAQ recovery procedure is still a big puzzle
- Coordination of the "experts on call" - e.g. TH needs improvement
- Beam trip rate is high and beam position moves a lot
Current Status
- Production on He-3 target when beam is available
Run Plan
- Continue running 45 uA on He3
- Accelerator will prioritize beam and not have any maintenance days this week
Data Collection and Polarization Status
- 37.7C of uncorrected charge for this kinematic (43.3C of corrected charge which is 28% of the plan)
- Last NMR 51.4%
Other Topics/Status Updates
- 8 Target Operator Shifts need to be taken
- Bill wants to change the raster settings after some analysis
Requested Accesses to the Hall
- Todd need to switch a gas bottle on GRINCH
- Jack would like to collect some cables